# Network Analytics for Anti-Money Laundering – A Systematic Literature Review and Experimental Evaluation

### v2

@@ Appendix B Analysis of Categories of Full Scope of Papers
Figure 1: The learning methods.
@@
Figure 2: The objective of the method.
@@
Figure 3: The evaluation metrics.
@@
Figure 4: The nature of the data.
@@ Appendix C Analysis of Categories between Crypto and Non-Crypto Papers
(a) The data for papers dealing with crypto currencies (1) and those that do not (0).
(b) The evolution of the data over the years for papers not covering crypto currencies.
(c) The evolution of the data over the years for papers covering crypto currencies.
Figure 1: Evolution of the data.
Figure 2: The distribution of the objective of the papers.
Figure 3: The distribution of the evaluation metrics used.
@@ Appendix D The Top-Cited Papers
Figure 4: The methods for the top-cited papers.
Figure 5: The performance metrics for the top-cited papers.
Figure 6: The objective of the top-cited papers.
Figure 7: The nature of the data of the top-cited papers.
@@ Appendix E Results Threshold-Dependent Metrics
Intrinsic features $0.6846\pm 0.1629$ $0.0135\pm 0.0034$ $0.0265\pm 0.0066$
Egonet features $0.6107\pm 0.1723$ $0.0120\pm 0.0035$ $0.0236\pm 0.0069$
Deepwalk $0.9989\pm 0.0133$ $0.0196\pm 0.0012$ $0.0384\pm 0.0023$
Node2vec $\boldsymbol{1.0000\pm 0.0000}$ $0.0196\pm 0.0012$ $0.0385\pm 0.0023$
GCN $\boldsymbol{1.0000\pm 0.0000}$ $\boldsymbol{0.2212\pm 0.0163}$ $\boldsymbol{0.3620\pm 0.0219}$
GraphSAGE $\boldsymbol{1.0000\pm 0.0000}$ $0.0314\pm 0.0000$ $0.0610\pm 0.0000$
GAT $0.9992\pm 0.0114$ $0.0198\pm 0.0018$ $0.0388\pm 0.0035$
GIN $0.9151\pm 0.0826$ $0.0175\pm 0.0019$ $0.0344\pm 0.0037$
Table 1: Threshold-dependent metrics: Precision, recall and F1-score values for the top $0.1\%$ scores over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.
@@
Intrinsic features $0.3449\pm 0.0113$ $0.6070\pm 0.0198$ $0.4398\pm 0.0144$
Egonet features $0.3276\pm 0.0105$ $0.5767\pm 0.0185$ $0.4179\pm 0.0134$
Deepwalk $0.3040\pm 0.0108$ $0.5352\pm 0.0191$ $0.3878\pm 0.0138$
Node2vec $0.3102\pm 0.0107$ $0.5460\pm 0.0188$ $0.3956\pm 0.0136$
GCN $0.3374\pm 0.0160$ $0.5941\pm 0.0187$ $0.4303\pm 0.0164$
GraphSAGE $\boldsymbol{0.3469\pm 0.0004}$ $\boldsymbol{0.6104\pm 0.0006}$ $\boldsymbol{0.4424\pm 0.0005}$
GAT $0.3270\pm 0.0169$ $0.5761\pm 0.0186$ $0.4170\pm 0.0172$
GIN $0.3216\pm 0.0162$ $0.5666\pm 0.0194$ $0.4102\pm 0.0169$
Table 2: Threshold-dependent metrics: Precision, recall and F1-score values for the top $10\%$ scores over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.
@@
Intrinsic features $0.3139\pm 0.0100$ $0.6103\pm 0.0194$ $0.4146\pm 0.0132$
Egonet features $0.2991\pm 0.0098$ $0.5815\pm 0.0190$ $0.3950\pm 0.0129$
Deepwalk $0.2775\pm 0.0098$ $0.5394\pm 0.0190$ $0.3664\pm 0.0129$
Node2vec $0.2850\pm 0.0095$ $0.5541\pm 0.0185$ $0.3764\pm 0.0126$
GCN $0.3109\pm 0.0145$ $0.6040\pm 0.0185$ $0.4104\pm 0.0154$
GraphSAGE $\boldsymbol{0.3147\pm 0.0000}$ $\boldsymbol{0.6116\pm 0.0000}$ $\boldsymbol{0.4156\pm 0.0000}$
GAT $0.2992\pm 0.0154$ $0.5817\pm 0.0186$ $0.3950\pm 0.0164$
GIN $0.2941\pm 0.0150$ $0.5718\pm 0.0197$ $0.3883\pm 0.0164$
Table 3: Threshold-dependent metrics: Precision, recall and F1-score values for the top $p=2\%$ scores over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.

### v3

@@ Appendix B Analysis of Categories of Full Scope of Papers
Figure B1: The learning methods.
(a) The distribution of the model building blocks.
(b) The evolution of the model building blocks over the years. Since there are many modelling methods, the graphs are split.
Figure B2: The building blocks of the model.
@@
Figure B3: The objective of the method.
@@
Figure B4: The evaluation metrics.
@@
Figure B5: The nature of the data.
@@ Appendix C Analysis of Categories between Crypto and Non-Crypto Papers
Figure C1: The data for papers dealing with crypto currencies (1) and those that do not (0).
(a) The evolution of the data over the years for papers not covering crypto currencies.
(b) The evolution of the data over the years for papers covering crypto currencies.
Figure C2: Evolution of the data.
Figure C3: The distribution of the objective of the papers.
Figure C4: The distribution of the evaluation metrics used.
@@ Appendix D The Top-Cited Papers
Figure D1: The methods for the top-cited papers.
Figure D2: The modelling methods for the top-cited papers.
Figure D3: The performance metrics for the top-cited papers.
Figure D4: The objective of the top-cited papers.
Figure D5: The nature of the data of the top-cited papers.
@@ Appendix E Results Threshold-Dependent Metrics
Table E1: Threshold-dependent metrics: Precision, recall and F1-score values for the top $0.1\%$ scores over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.
@@
Intrinsic features $0.061\pm 0.117$ $0.001\pm 0.002$ $0.002\pm 0.005$
Egonet features $0.589\pm 0.186$ $0.012\pm 0.004$ $0.023\pm 0.007$
DeepWalk $0.637\pm 0.147$ $0.013\pm 0.003$ $0.025\pm 0.006$
Node2vec $0.984\pm 0.048$ $0.019\pm 0.002$ $0.038\pm 0.003$
GCN $\boldsymbol{1.000\pm 0.000}$ $0.019\pm 0.001$ $0.037\pm 0.001$
GraphSAGE $0.942\pm 0.073$ $\boldsymbol{0.020\pm 0.003}$ $\boldsymbol{0.039\pm 0.005}$
GAT $\boldsymbol{1.000\pm 0.000}$ $0.019\pm 0.001$ $0.037\pm 0.001$
GIN $\boldsymbol{1.000\pm 0.000}$ $0.019\pm 0.001$ $0.038\pm 0.002$
Table E2: Threshold-dependent metrics: Precision, recall and F1-score values for the top $0.1\%$ scores over the different methods for the IBM-AML data set, based on the test set. The standard deviation is also reported.
@@
Intrinsic features $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
Egonet features $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
DeepWalk $\boldsymbol{0.018\pm 0.012}$ $0.002\pm 0.002$ $0.004\pm 0.003$
Node2vec $0.007\pm 0.009$ $0.001\pm 0.001$ $0.002\pm 0.002$
GCN $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
GraphSAGE $0.009\pm 0.000$ $\boldsymbol{1.000\pm 0.000}$ $\boldsymbol{0.019\pm 0.001}$
GAT $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
GIN $0.010\pm 0.010$ $0.001\pm 0.001$ $0.002\pm 0.002$
Table E3: Threshold-dependent metrics: Precision, recall and F1-score values for the top $10\%$ scores over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.
@@
Intrinsic features $0.357\pm 0.011$ $0.629\pm 0.019$ $0.456\pm 0.014$
Egonet features $0.361\pm 0.010$ $0.636\pm 0.018$ $0.460\pm 0.013$
DeepWalk $0.347\pm 0.011$ $0.611\pm 0.019$ $0.443\pm 0.014$
Node2vec $0.351\pm 0.011$ $0.617\pm 0.019$ $0.447\pm 0.014$
GCN $0.348\pm 0.019$ $0.614\pm 0.021$ $0.444\pm 0.020$
GraphSAGE $0.358\pm 0.016$ $0.633\pm 0.018$ $0.458\pm 0.016$
GAT $\boldsymbol{0.379\pm 0.015}$ $\boldsymbol{0.665\pm 0.018}$ $\boldsymbol{0.482\pm 0.015}$
GIN $0.341\pm 0.016$ $0.600\pm 0.019$ $0.435\pm 0.016$
Table E4: Threshold-dependent metrics: Precision, recall and F1-score values for the top $10\%$ scores over the different methods for the IBM-AML data set, based on the test set. The standard deviation is also reported.
Methods Precision Recall F1-score
Intrinsic features $0.015\pm 0.001$ $0.161\pm 0.013$ $0.027\pm 0.002$
Egonet features $0.014\pm 0.001$ $0.157\pm 0.012$ $0.026\pm 0.002$
DeepWalk $\boldsymbol{0.015\pm 0.001}$ $0.165\pm 0.012$ $\boldsymbol{0.028\pm 0.002}$
Node2vec $0.015\pm 0.001$ $0.164\pm 0.013$ $0.027\pm 0.002$
GCN $0.012\pm 0.000$ $0.986\pm 0.004$ $0.023\pm 0.001$
GraphSAGE $0.009\pm 0.000$ $\boldsymbol{1.000\pm 0.000}$ $0.019\pm 0.001$
GAT $0.002\pm 0.000$ $0.017\pm 0.004$ $0.003\pm 0.001$
GIN $0.015\pm 0.001$ $0.160\pm 0.012$ $0.027\pm 0.002$
Table E5: Threshold-dependent metrics: Precision, recall and F1-score values for the top $p=2\%$ scores over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.
Methods Precision Recall F1-score
Intrinsic features $0.327\pm 0.010$ $0.635\pm 0.019$ $0.432\pm 0.013$
Egonet features $0.334\pm 0.009$ $0.650\pm 0.018$ $0.441\pm 0.012$
DeepWalk $0.316\pm 0.010$ $0.614\pm 0.019$ $0.417\pm 0.013$
Node2vec $0.318\pm 0.010$ $0.619\pm 0.019$ $0.420\pm 0.013$
GCN $0.317\pm 0.017$ $0.618\pm 0.020$ $0.419\pm 0.019$
GraphSAGE $0.329\pm 0.015$ $0.641\pm 0.018$ $0.435\pm 0.015$
GAT $\boldsymbol{0.350\pm 0.014}$ $\boldsymbol{0.677\pm 0.018}$ $\boldsymbol{0.461\pm 0.015}$
GIN $0.315\pm 0.015$ $0.612\pm 0.020$ $0.416\pm 0.016$
Table E6: Threshold-dependent metrics: Precision, recall and F1-score values for the top $p=0.11\%$ scores over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.
Methods Precision Recall F1-score
Intrinsic features $0.001\pm 0.004$ $0.000\pm 0.001$ $0.000\pm 0.001$
Egonet features $0.001\pm 0.003$ $0.000\pm 0.001$ $0.000\pm 0.001$
DeepWalk $\boldsymbol{0.017\pm 0.012}$ $0.002\pm 0.002$ $0.004\pm 0.003$
Node2vec $0.011\pm 0.009$ $0.002\pm 0.001$ $0.003\pm 0.003$
GCN $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
GraphSAGE $0.009\pm 0.000$ $\boldsymbol{1.000\pm 0.000}$ $\boldsymbol{0.019\pm 0.001}$
GAT $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000
GIN $0.009\pm 0.009$ $0.001\pm 0.001$ $0.002\pm 0.002$